# Teaching Human Behavior Improves Content Understanding Abilities Of LLMs

### v1

@@ 2.2 Instruction Fine-Tuning LLaMA-Vid
Listing 2: Perceptual Signal Instruction fine-tuning template for the image: http://farm6.staticflickr.com/5106/5670500150_e035dd2d30_z.jpg
@@
As an ablation experiment, we also try teaching the Behavior-LLaVA perceptual signals. For this, we take the largest perception signal dataset in the literature - Salicon10k [36]. It consists of 10,000 MS COCO images [18] with free-viewing eye gaze data collected through a novel mouse-based interface. The dataset has been widely used in many studies. We formulate two tasks using this data, (1) Salicon [Object]: estimating the saliency over the objects in the image and (2) Salicon [Region]: estimating the saliency over a region, where the regions tiles formed by breaking the image into a 3x3 grid. For both tasks we try to model two objectives, ranking and predicting, we found ranking to be much more effective. The instruction is given in Listings 2 and 3.
Listing 3: Perceptual Signal Instruction fine-tuning template for the image: http://farm6.staticflickr.com/5106/5670500150_e035dd2d30_z.jpg
@@
We take the pretrained checkpoints of the LLaMA-VID 13B [51] model after the modality-alignment stage. We combine our instruction pairs from BLIFT with the original instruction tuning dataset consisting of 40K text conversations from ShareGPT, 625K single or multi-turn visual QA pairs, and 98K video QA pairs, all the modules except the Visual Encoder are kept frozen. We train the LLaMA-Vid checkpoints with their original sft mix along with BLIFT. We ablate different sampling ratios and found 1:1 to be empirically performing the best. We train the model for 2.2 epochs, keeping track of the 0-shot evaluation metrics and perplexity on comments in the eval set. For the best checkpoint, the perplexity on comments reduces from 6.22 to 3.05, and the $R^{2}$ on likes/views goes from -5.1 to 0.45.
3 Results and Discussion
Model scene way_speaking relationship like_ratio view_count director genre writer year Video4096-GPT-3.5 generated story + Flan-t5-xxl [13] 60.2 39.07 64.1 0.061 12.84 69.9 58.1 52.4 75.6 Video4096-GPT-3.5 generated story + GPT-3.5 classifier [13] 54.54 32.95 68.42 0.031 12.69 75.26 50.84 32.16 75.96 LLaMA-Vid + GPT-3.5 Generated Story 58.12 35.5 60.6 0.314 10.34 65.34 49.77 34.23 72.12 Ad-LLaVA 59.05 37.07 61.2 0.319 10.37 66.84 55.13 35.33 77.34 Behavior LLaVA + GPT-3.5 Generated Story 66.43 41.03 64.21 0.17 5.12 71.12 63.45 39.4 79.3 Improvement of Behavior LLaVA over LLaMA-Vid 10.48% 15.58% 9.62% 45.86% 50.48% 8.85% 27.49% 15.1% 9.96%
Table 1: Comparison of various models on the Long Video Understanding benchmark [79] consisting of 9 VQA tasks. We see that Behavior-LLaVA improves on LLaMA-Vid on 9/9 tasks with an average improvement of 21.49%. Further, it outperforms the state-of-the-art in 5/9 tasks.
Model Scene Object Action Event Attribute Concept Overall 0-shot LLaMA-VID 47.25 65.26 78.12 28.03 42.33 50.03 51.83 Ad-LLaVA 49.10 65.35 77.45 31.45 43.33 50.70 52.91 Behavior-LLaVA 52.03 65.33 77.95 32.66 45.67 51.20 54.14 Improvement of Behavior-LLaVA over LLaMA-Vid 10.12% 0.11% -0.22% 16.52% 7.89% 2.34% 4.46% 0-shot w/ story Video4096 - GPT-3.5 generated story + Flan-t5-xxl classifier [13] 59.66 98.89 98.96 38.42 67.76 86.99 75.12 Video4096 - GPT-3.5 generated story + GPT-3.5 classifier [13] 60.2 99.16 98.72 40.79 67.17 88.6 75.77 LLaMA-VID + Generated Story 60.3 99.92 99.01 39.33 66.66 87.33 75.425 Behavior-LLaVA + GPT3.5 generated story 60.4 99.89 98.23 40.97 67.23 88.33 75.84 Improvement of Behavior-LLaVA over LLaMA-Vid 0.16% -0.03% -0.78% 4.17% 0.85% 1.14% 0.55%
Table 2: Comparison of various models on the Holistic Video Understanding benchmark [23] consisting of 7 VQA tasks. We see that Behavior-LLaVA improves on LLaMA-Vid on 6/7 tasks with an average improvement of 5.88%. Further, it performs equivalent to the state-of-the-art in 6/7 tasks. State of the art is achieved by generating a story and asking Behavior-LLaVA to answer questions based on the generated story.
3.1 Evaluation
To test the effectiveness of Behavior-LLaVA, we conduct experiments involving 40 distinct tasks across 23 benchmark datasets. The diversity of tasks and datasets allows us to evaluate the performance and capabilities of Behavior-LLaVA thoroughly. Each of them is covered briefly next:
Method LLM MSVD-QA MSRVTT-QA ActivityNet-QA Acc Score Acc Score Acc Score FrozenBiLM [84] DeBERTa-V2 32.2 16.8 24.7 VideoLLaMA [87] Vicuna-7B 51.6 2.5 29.6 1.8 12.4 1.1 LLaMA-Adapter [88] LLaMA-7B 54.9 3.1 43.8 2.7 34.2 2.7 VideoChat [50] Vicuna-7B 56.3 2.8 45.0 2.5 26.5 2.2 Video-ChatGPT [56] Vicuna-7B 64.9 3.3 49.3 2.8 35.2 2.7 BT-Adapter [55] Vicuna-7B 67.5 3.7 57.0 3.2 45.7 3.2 LLaMA-VID Vicuna-7B 69.7 3.7 57.7 3.2 47.4 3.3 LLaMA-VID Vicuna-13B 70.0 3.7 58.9 3.3 47.5 3.3 Ad-LLaVA Vicuna-13B 70.0 3.7 59.0 3.3 47.4 3.3 Behaviour-LLaVA Vicuna-13B 70.1 3.7 59.2 3.4 47.5 3.3 Improvement of Behavior-LLaVA over LLaMA-Vid 0.14% 0% 0.5% 3% 0% 0%
  1. 1.
    • The Long-Video Understanding (LVU) benchmark by Wu and Krahenbuhl [79] comprises nine distinct tasks aimed at assessing long video comprehension, incorporating over 1000 hours of video content. These tasks encompass diverse aspects such as content understanding (including ‘relationship’, ‘speaking style’, ‘scene/place’), prediction of user engagement (‘YouTube like ratio’, ‘YouTube popularity’), and forecasting movie metadata (‘director’, ‘genre’, ‘writer’, ‘movie release year’). Evaluation of LVU tasks by Wu and Krahenbuhl [79] utilizes top-1 classification accuracy for content understanding and metadata prediction tasks, while Mean Squared Error (MSE) is employed for user engagement prediction tasks.
    • The Holistic Video Understanding (HVU) dataset, as introduced by Diba et al. [23], stands as the largest dataset for long video comprehension, comprising 572,000 samples. Encompassing a broad spectrum of semantic elements within videos, HVU tasks involve the classification of scenes, objects, actions, events, attributes, and concepts. Performance evaluation on HVU tasks, akin to the original work, is conducted using the mean average precision (mAP) metric on the validation set.
    • MSVD-QA and MSRVTT-QA: These datasets are based on Microsoft Research Video Description [17] and MSR-VTT corpora [82] and are extensively used in many video captioning and question-answering experiments. The MSVD-QA dataset has a total number of 1,970 video clips and 50,505 question-answer pairs. The MSRVTT-QA dataset contains 10K video clips and 243k question-answer pairs.
    • ActivityNet-QA [16] is a benchmark primarily for human activity understanding containing 849 hours of video, including 28,000 action instances.
  2. 2.
    • The advertisements dataset by Hussain et al. [33] contains 3,477 video advertisements and the corresponding annotations for emotion and topic tags and action-reason statements for each video. There are a total of 38 topics and 30 unique emotion tags per video. Further, we have 5 action-reason statements for each video for the action-reason generation task.
    • Persuasion strategy dataset [13] is a dataset consisting of 1002 video advertisements from popular brands and their persuasion strategy labels like social identity, anchoring and comparison, reciprocity, foot-in-the-door, etc.
      Training Model Topic Sentiment Persuasion Action Reason Clubbed All labels Random Random 2.63 3.37 14.3 8.37 3.34 3.33 Zero-shot VideoChat [50] 9.07 3.09 5.1 10.28 - - Video4096 - GPT-3.5 Generated Story + GPT-3.5 Classifier [13] 51.6 11.68 79.69 35.02 66.27 59.59 Video4096 - GPT-3.5 Generated Story + Flan-t5-xxl Classifier [13] 60.5 10.8 79.10 33.41 79.22 81.72 Video4096 - GPT-3.5 Generated Story + Vicuna Classifier [13] 22.92 10.8 67.35 29.6 21.39 20.89 Video4096 - Vicuna Generated Story + GPT-3.5 Classifier [13] 46.7 5.9 80.33 27.54 61.88 55.44 Video4096 - Vicuna Generated Story + Flan-t5-xxl Classifier [13] 57.38 9.8 76.60 30.11 77.38 80.66 Video4096 - Vicuna Generated Story + Vicuna Classifier [13] 11.75 10.5 68.13 26.59 20.72 21.00 LCBM [39] 42.17 7.08 58.83 32.83 39.55 27.91 LLaMA-VID w/ only video 10.11 3.42 5.75 12.32 29.61 24.11 LLaMA-VID w/ video + GPT-3.5 Story 42.72 11.05 64.02 32.07 37.76 42.33 Behavior-LLaVA w/ only video 22.65 11.13 60.04 13.39 42.66 33.33 Behavior-LLaVA w/ video + verbalization 46.34 11.7 64.13 33.33 52.06 52.03 Ad-LLaVA w/ video + GPT-3.5 story 51.16 11.33 68.03 33.11 43.26 51.45 Behavior-LLaVA w/ video + GPT-3.5 story 60.09 12.84 79.94 36.12 67.10 79.18 Improvement of Behavior-LLaVA over LLaMA-Vid 40.66% 16.2% 24.86% 12.62% 77.7% 87.05% Finetuned VideoMAE [75] 24.72 29.72 85.55 11.17 - - Hussain et al. [33] 35.1 32.8 - - 48.45 - Intern-Video [78] 57.47 36.08 86.59 5.47 6.8 - Video4096- Generated Story + Roberta Classifier [13] 71.3 33.02 84.20 64.67 42.96 39.09 LLaMA-VID w/ video + verbalization 59.13 32.11 79.15 50.93 50.32 30.13 LLaMA-VID w/ video + GPT-3.5 Story 63.11 35.01 84.15 55.01 57.11 45.73 Behavior-LLaVA w/ only video 58.03 22.72 84.41 26.23 59.33 51.45 Behavior-LLaVA w/ video + verbalization 68.32 33.92 85.93 64.72 70.89 75.34 Ad-LLaVA w/ video + GPT-3.5 story 66.34 36.24 84.09 58.31 68.15 78.15 Behavior-LLaVA w/ video + GPT-3.5 story 71.2 39.55 86.17 65.03 80.44 81.67 Improvement of Behavior-LLaVA over LLaMA-Vid 12.82% 12.97% 2.4% 18.21% 40.85% 78.59%
      Table 4: Comparison of various models on two video understanding benchmarks [33, 44] consisting of 5 tasks related to video advertisements understanding. The main goal of comparing on these benchmarks is to demonstrate Behavior-LLaVA’s understanding of complex videos. We see that Behavior-LLaVA improves on LLaMA-Vid on 5/5 tasks with an average improvement score of 43.18% in zero-shot and 27.64% in fine-tuned settings. Further, it outperforms the current state-of-the-art on 3/5 tasks in zero-shot and 5/5 in fine-tuned settings.
    • VideoEmotion-8 [7] dataset comprises 1,101 user-generated videos sourced from YouTube and Flickr, each containing a minimum of 100 videos per emotional category, as per Plutchik Wheel’s emotion model.
    • Ekman-6 [81] dataset is compiled from social websites, with each of its 1,637 videos labeled with a single emotion category based on Ekman’s psychological research.
    • CAER [47] dataset, sourced from TV shows, consists of 13,201 clips with an average sequence length of 90, each manually labeled with six basic emotions, aligning with the Ekman-6 dataset.
    • IAPSa [60] is a subset of IAPS, following the Mikels model with eight emotion categories such as amusement, awe, contentment, excitement, anger, disgust, fear, and sadness. It consists of 395 affective images, marking the first visual emotion dataset with discrete categories.
    • Emotion6 [64] features 1,980 images sourced from Flickr, each labeled by 15 annotators according to the Ekman model, covering six emotion categories: happiness, anger, disgust, fear, sadness, and surprise.
    • EmoSet [86] encompasses a total of 3.3 million images, including 118,102 from social networks and artistic sources, evenly distributed across various emotion categories. Based on the Mikels model, EmoSet is categorized into eight emotion categories.
    • Abstract [57] exclusively consists of color and texture combinations without recognizable objects. Differing from the IAPS dataset where emotions often stem from identifiable objects, the abstract paintings dataset was peer-rated via a web survey, with each image rated approximately 14 times. It comprises 228 images spanning eight categories similar to those in IAPS.
  3. 3.
    Image Dense Captioning: Literature image captioning datasets such as MS-COCO [18] reduce the inherently rich information and fine-grained semantics to simplistic captions, with very brief statements often focussing only on salient objects. This leads to a significant reduction in information content. Behavior data such as user comments help a model learn much more information such as object and material properties, world knowledge, emotion, character understanding, spatial relationships, aesthetics, etc. (see Fig. 1), enhancing the model’s captioning capability. Therefore, we design a captioning task to test this capability and compare it with respect to LLaMA-Vid and LLaVA-34B (a 2.5x larger model). Since we do not have ground truth for this task, following the LLM-as-a-judge paradigm, we use GPT-4V as the judge for all the models. GPT-4V is asked to evaluate the dense captions on three metrics: Correctness (Listing 4) evaluating the factuality and model hallucinations, Detail (Listing 5) evaluating the number and depth of details captured by the generated captions, and Quality (Listing 6) measuring the subjective quality of the concepts chosen to be highlighted by the captioning model and the arrangement, coherence, and the linking of various concepts.
  4. 4.
    Image and Video Memorability Simulation: Behavior-LLaVA is trained on behavior along with the media. To check if training on behavior helps in solving other behavior tasks [39], we test it over image and video memorability simulation. For this, we select seven benchmark datasets covering long-term and short-term memorability over images and videos.
    • Memento10k[61] is a short-term video memorability dataset comprising 10,000 video clips, with 900,000 human memory annotations recorded at various delay intervals. The video clips were, on average, 3s long.
    • VideoMem [20] is comprised of 10,000 soundless videos, each lasting 7 seconds, accompanied by memorability scores. Memorability was measured twice: first, shortly after viewing and again 24-72 hours later to capture both short-term and long-term memorability effects.
    • LaMem [40] dataset is a short-term image memorability dataset comprising of 60000 images. The dataset contains scene-centric images, object-centric images and other types such as images of art, images evoking certain emotions, and other user-generated images.
    • SUN [35] dataset is a short-term image memorability dataset comprising $2$ images that were sourced from the SUN database.
      Training Dataset Video Emotion-8 CAER Ekman-6 Random Random 12.5 14.28 16.67 0-Shot LLaMA-Vid 29.7 27.2 37.33 Behavior-LLaVA 41.35 51.0 49.33 Ad-LLaVA 29.8 27.3 37.66 Improvement of Behavior-LLaVA over LLaMA-Vid 39.22% 84.19% 32.14% Finetuned Zhao et al. [90] 54.5 78.3 55.3 Zhang et al. [89] 57.3 80.1 58.2 eMOTIONS [80] - - 53.12 Arevalo et al. [6] 53.7 77.3 54.2 Qiu et al. [68] 53.3 - 57.3 Xu et al. [81] 52.6 77.9 55.6 LLaMA-Vid 53.8 75.6 57.9 Ad-LLaVA 54.1 76.1 57.8 Behavior-LLaVA 56.9 79.3 58.4 Improvement of Behavior-LLaVA over LLaMA-Vid 5.76% 3.57% 0.86%
      Table 5: Comparison of various models on three video emotion understanding benchmarks (Video Emotion8 [37], CAER [47], Ekman-6 [81]). The main goal of comparing on these benchmarks is to demonstrate Behavior-LLaVA’s understanding of complex tasks like video emotions of long-form videos. We see that Behavior-LLaVA improves on LLaMA-Vid on 3/3 benchmarks with an average improvement score of 51.85% in zero-shot and 3.39% in fine-tuned settings. Further, it outperforms the current state-of-the-art on 3/3 benchmarks in zero-shot and 1/3 in fine-tuned settings.
    • MemCat [27] dataset is a short-term image memorability dataset comprising 10000 images. It consists of five broader memorability-relevant semantic categories (animal, sports, food, landscapes, vehicles), with 2K exemplars each, further divided into different subcategories (e.g., bear, pigeon, cat, etc. for animal). The images were sourced from existing image sets: ImageNet, COCO, Open Images Dataset, and SUN.
    • MediaEval [42] utilized publicly available links to short-form video clips, each averaging 6 seconds in duration, with both short-term and long-term memorability scores. Short-term memorability evaluations were conducted on videos viewed within the preceding few minutes, while long-term memorability assessments were based on videos viewed within the previous 24 to 72 hours.
    • LAMBDA [34] is a long-term memorability consisting of 2205 video ads collected over 1749 participants covering 276 brands. The average video length is 33 seconds and the videos are highly complex, consisting of audio, logos, fast-moving scenes, emotions, etc.
For Tables 1, 2, and 4, we follow the evaluation protocol of Video-4096 [13], for Table 3 we follow the evaluation protocol of LLaVA and LLaMA-VID. For Tables 5, 6, and 8, for 0-shot evaluation results, we use the logits of the next token from the given task vocabulary. For Table 8, we use the evaluation protocol of SI et al. [73].
@@ 3.1 Evaluation
Training Models IAPSa-8 Abstract Emotion6 Emoset Random Random 12.5 12.5 16.67 12.5 0-shot GPT4-V 83.33 71.12 65.47 79.16 LLaMA-Vid 43.41 43.24 40.37 45.23 Ad-LLaVA 43.22 43.01 43.21 44.38 Behavior-LLaVA 57.97 64.21 49.71 50.38 Improvement of Behavior-LLaVA over LLaMA-Vid 33.54% 48.5% 23.14% 11.39% Finetuned MIDAN [83] 85.96 78.34 61.66 75.75 Stimuli-aware [85] - - 61.62 78.40 LLaMA-VID finetuned 84.93 71.23 62.87 80.31 Ad-LLaVA finetuned 85.13 71.16 62.66 79.88 Behavior-LLaVA finetuned 87.36 81.41 72.31 83.21 Improvement of Behavior-LLaVA over LLaMA-Vid 2.86% 14.29% 15.02% 3.61%
Table 6: Comparison of various models on four image emotion understanding benchmarks (IAPSa-8 [60] Abstract [57], Emotion6 [64], Emoset [86]). The main goal of comparing on these benchmarks is to demonstrate Behavior-LLaVA’s understanding of complex tasks like image emotions. We see that Behavior-LLaVA improves on LLaMA-Vid on 4/4 benchmarks with an average improvement score of 29.14% in zero-shot and 8.95% in fine-tuned settings. Further, it outperforms the current state-of-the-art on 4/4 benchmarks in the fine-tuned settings.
@@
Training Models Image Datasets Video Datasets Lamem Memcat SUN Memento10k VideoMem MediaEval LAMBDA Human Consistency 0.68 0.78 0.75 0.73 0.61 - 0.61 Finetuned 10-shot in-context learning GPT-3.5 0.29 0.18 0.15 0.07 0.06 0.06 0.06 ViTMem [28] 0.71 0.65 0.63 0.56 0.51 - 0.08 Henry trained with 25% data [34] 0.56 0.64 0.59 0.62 0.49 0.32 0.28 Henry trained with 50% data [34] 0.65 0.68 0.67 0.69 0.55 0.44 0.40 Henry trained with 75% data [34] 0.71 0.75 0.73 0.74 0.62 0.49 0.47 Henry trained on all (combined) datasets [34] 0.72 0.79 0.76 0.72 0.60 0.48 0.52 Ad-LLaVA trained with 50% data 0.67 0.65 0.61 0.69 0.56 0.43 0.47 Behaviour LLaVA trained with 25% data 0.67 0.72 0.69 0.68 0.53 0.44 0.50 Behaviour LLaVA trained with 50% data 0.72 0.77 0.73 0.71 0.59 0.46 0.51 Behaviour LLaVA trained with 75% data 0.73 0.77 0.74 0.70 0.60 0.47 0.50 Behavior-LLaVA trained on all datasets 0.73 0.78 0.74 0.71 0.60 0.47 0.52 Improvement of Behavior-LLaVA over LLaMA-Vid (25% data) 19.64% 12.5% 16.95% 9.68% 8.16% 37.5% 78.57% Improvement of Behavior-LLaVA over LLaMA-Vid (50% data) 10.77% 13.26% 8.96% 2.90% 7.27% 4.54% 27.5% 0-shot LLaMA-Vid 0.13 0.11 0.05 0.03 0.05 0.02 0.05 Ad-LLaVA 0.14 0.13 0.06 0.06 0.07 0.04 0.13 Behavior-LLaVA 0.21 0.17 0.13 0.12 0.08 0.07 0.16 Improvement of Behavior-LLaVA over LLaMA-Vid 61.5% 54.5% 160% 300% 160% 350% 219%
Table 8: Comparison of various models on seven video and image memorability benchmarks (Memento10k [61], VideoMem [20], LaMem [40], SUN [35], MemCat [27], MediaEval [42], LAMBDA [34]). The main goal of comparing on these benchmarks is to demonstrate Behavior-LLaVA’s understanding of complex and high-level tasks like memorability simulation. We see that Behavior-LLaVA improves on LLaMA-Vid on 7/7 benchmarks with an average improvement score of 186.4% in zero-shot and 39% in fine-tuned settings after seeing 25% train data. Further, it performs similarly to the current state-of-the-art on 7/7 benchmarks in the fine-tuned settings while still seeing only 25% data.
Task MSRVTT-QA Emoset Comments likes/views Base 58.9 45.23 6.22 -0.1 Salicon [Region] 55.6 47.25 6.25 -0.07 Salicon [Object] 57.9 48.0 6.12 0.05 BLIFT[Likes/Views] 58.2 47.12 6.15 0.38 BLIFT[Comments] 59.0 49.58 3.02 0.19 BLIFT 59.2 50.38 3.05 0.40

### v3

@@ 4 Conclusion
Listing 5: Perceptual Signal Instruction fine-tuning template for the image: http://farm6.staticflickr.com/5106/5670500150_e035dd2d30_z.jpg
@@
Listing 6: Perceptual Signal Instruction fine-tuning template for the image: http://farm6.staticflickr.com/5106/5670500150_e035dd2d30_z.jpg
@@
Figure 5: Dense caption generated by Behavior-LLaVA for the video of a Volkswagen ad. The original video is posted at URL: https://www.youtube.com/watch?v=kyuGXPNr-T0. The red-colored text highlights the most important aspects of the video captured by Behavior-LLaVA, demonstrating an understanding of aesthetics, characters, world knowledge, emotion, and spatial relationships. More such examples are given in Figs. 6, 7, 10, and Figs. 8, 9 for images and videos respectively.
@@
Training Model Topic Sentiment Persuasion Action Reason Clubbed All labels Random Random 2.63 3.37 14.3 8.37 3.34 3.33 Zero-shot VideoChat (Li et al., 2023a) 9.07 3.09 5.1 10.28 - - Video4096 - GPT-3.5 Generated Story + GPT-3.5 Classifier (Bhattacharyya et al., 2023) 51.6 11.68 79.69 35.02 66.27 59.59 Video4096 - GPT-3.5 Generated Story + Flan-t5-xxl Classifier (Bhattacharyya et al., 2023) 60.5 10.8 79.10 33.41 79.22 81.72 Video4096 - GPT-3.5 Generated Story + Vicuna Classifier (Bhattacharyya et al., 2023) 22.92 10.8 67.35 29.6 21.39 20.89 Video4096 - Vicuna Generated Story + GPT-3.5 Classifier (Bhattacharyya et al., 2023) 46.7 5.9 80.33 27.54 61.88 55.44 Video4096 - Vicuna Generated Story + Flan-t5-xxl Classifier (Bhattacharyya et al., 2023) 57.38 9.8 76.60 30.11 77.38 80.66 Video4096 - Vicuna Generated Story + Vicuna Classifier (Bhattacharyya et al., 2023) 11.75 10.5 68.13 26.59 20.72 21.00 LCBM (Khandelwal et al., 2024) 42.17 7.08 58.83 32.83 39.55 27.91 LLaMA-VID w/ only video 10.11 3.42 5.75 12.32 29.61 24.11 LLaMA-VID w/ video + GPT-3.5 Story 42.72 11.05 64.02 32.07 37.76 42.33 Behavior-LLaVA w/ only video 22.65 11.13 60.04 13.39 42.66 33.33 Behavior-LLaVA w/ video + verbalization 46.34 11.7 64.13 33.33 52.06 52.03 Ad-LLaVA w/ video + GPT-3.5 story 51.16 11.33 68.03 33.11 43.26 51.45 Behavior-LLaVA w/ video + GPT-3.5 story 60.09 12.84 79.94 36.12 67.10 79.18 Improvement of Behavior-LLaVA over LLaMA-Vid 40.66% 16.2% 24.86% 12.62% 77.7% 87.05% Finetuned VideoMAE (Tong et al., 2022) 24.72 29.72 85.55 11.17 - - Hussain et al. (2017) 35.1 32.8 - - 48.45 - Intern-Video (Wang et al., 2022) 57.47 36.08 86.59 5.47 6.8 - Video4096- Generated Story + Roberta Classifier (Bhattacharyya et al., 2023) 71.3 33.02 84.20 64.67 42.96 39.09 LLaMA-VID w/ video + verbalization 59.13 32.11 79.15 50.93 50.32 30.13 LLaMA-VID w/ video + GPT-3.5 Story 63.11 35.01 84.15 55.01 57.11 45.73 Behavior-LLaVA w/ only video 58.03 22.72 84.41 26.23 59.33 51.45 Behavior-LLaVA w/ video + verbalization 68.32 33.92 85.93 64.72 70.89 75.34 Ad-LLaVA w/ video + GPT-3.5 story 66.34 36.24 84.09 58.31 68.15 78.15 Behavior-LLaVA w/ video + GPT-3.5 story 71.2 39.55 86.17 65.03 80.44 81.67 Improvement of Behavior-LLaVA over LLaMA-Vid 12.82% 12.97% 2.4% 18.21% 40.85% 78.59%
Table 6: Comparison of various models on two video understanding benchmarks (Hussain et al., 2017; Kumar et al., 2023) consisting of 5 tasks related to video advertisements understanding. The main goal of comparing on these benchmarks is to demonstrate Behavior-LLaVA’s understanding of complex videos. We see that Behavior-LLaVA improves on LLaMA-Vid on 5/5 tasks with an average improvement score of 43.18% in zero-shot and 27.64% in fine-tuned settings. Further, it outperforms the current state-of-the-art on 3/5 tasks in zero-shot and 5/5 in fine-tuned settings.
Method LLM MSVD-QA MSRVTT-QA ActivityNet-QA Acc Score Acc Score Acc Score FrozenBiLM (Yang et al., 2022) DeBERTa-V2 32.2 16.8 24.7 VideoLLaMA (Zhang et al., 2023a) Vicuna-7B 51.6 2.5 29.6 1.8 12.4 1.1 LLaMA-Adapter (Zhang et al., 2023b) LLaMA-7B 54.9 3.1 43.8 2.7 34.2 2.7 VideoChat (Li et al., 2023a) Vicuna-7B 56.3 2.8 45.0 2.5 26.5 2.2 Video-ChatGPT (Maaz et al., 2023) Vicuna-7B 64.9 3.3 49.3 2.8 35.2 2.7 BT-Adapter (Liu et al., 2023b) Vicuna-7B 67.5 3.7 57.0 3.2 45.7 3.2 LLaMA-VID Vicuna-7B 69.7 3.7 57.7 3.2 47.4 3.3 LLaMA-VID Vicuna-13B 70.0 3.7 58.9 3.3 47.5 3.3 Ad-LLaVA Vicuna-13B 70.0 3.7 59.0 3.3 47.4 3.3 Behaviour-LLaVA Vicuna-13B 70.1 3.7 59.2 3.4 47.5 3.3 Improvement of Behavior-LLaVA over LLaMA-Vid 0.14% 0% 0.5% 3% 0% 0%
Training Models IAPSa-8 Abstract Emotion6 Emoset Random Random 12.5 12.5 16.67 12.5 0-shot GPT4-V 83.33 71.12 65.47 79.16 LLaMA-Vid 43.41 43.24 40.37 45.23 Ad-LLaVA 43.22 43.01 43.21 44.38 Behavior-LLaVA 57.97 64.21 49.71 50.38 Improvement of Behavior-LLaVA over LLaMA-Vid 33.54% 48.5% 23.14% 11.39% Finetuned MIDAN (Xu et al., 2022) 85.96 78.34 61.66 75.75 Stimuli-aware (Yang et al., 2021) - - 61.62 78.40 LLaMA-VID finetuned 84.93 71.23 62.87 80.31 Ad-LLaVA finetuned 85.13 71.16 62.66 79.88 Behavior-LLaVA finetuned 87.36 81.41 72.31 83.21 Improvement of Behavior-LLaVA over LLaMA-Vid 2.86% 14.29% 15.02% 3.61%
Table 8: Comparison of various models on four image emotion understanding benchmarks (IAPSa-8 (Mikels et al., 2005) Abstract (Machajdik & Hanbury, 2010), Emotion6 (Peng et al., 2015), Emoset (Yang et al., 2023)). The main goal of comparing on these benchmarks is to demonstrate Behavior-LLaVA’s understanding of complex tasks like image emotions. We see that Behavior-LLaVA improves on LLaMA-Vid on 4/4 benchmarks with an average improvement score of 29.14% in zero-shot and 8.95% in fine-tuned settings. Further, it outperforms the current state-of-the-art on 4/4 benchmarks in the fine-tuned settings.
@@
Model Scene Object Action Event Attribute Concept Overall 0-shot LLaMA-VID 47.25 65.26 78.12 28.03 42.33 50.03 51.83 Ad-LLaVA 49.10 65.35 77.45 31.45 43.33 50.70 52.91 Behavior-LLaVA 52.03 65.33 77.95 32.66 45.67 51.20 54.14 Improvement of Behavior-LLaVA over LLaMA-Vid 10.12% 0.11% -0.22% 16.52% 7.89% 2.34% 4.46% 0-shot w/ story Video4096 - GPT-3.5 generated story + Flan-t5-xxl classifier 59.66 98.89 98.96 38.42 67.76 86.99 75.12 Video4096 - GPT-3.5 generated story + GPT-3.5 classifier 60.2 99.16 98.72 40.79 67.17 88.6 75.77 LLaMA-VID + Generated Story 60.3 99.92 99.01 39.33 66.66 87.33 75.425 Behavior-LLaVA + GPT3.5 generated story 60.4 99.89 98.23 40.97 67.23 88.33 75.84 Improvement of Behavior-LLaVA over LLaMA-Vid 0.16% -0.03% -0.78% 4.17% 0.85% 1.14% 0.55%
Table 10: Comparison of various models on the Holistic Video Understanding benchmark (Diba et al., 2020) consisting of 7 VQA tasks. We see that Behavior-LLaVA improves on LLaMA-Vid on 6/7 tasks with an average improvement of 5.88%. Further, it performs equivalent to the state-of-the-art in 6/7 tasks. State of the art is achieved by generating a story and asking Behavior-LLaVA to answer questions based on the generated story.
Task MSRVTT-QA CAER Emoset Comments likes/views Base 58.9 75.6 45.23 6.22 -0.1 Salicon [Region] 55.6 75.8 47.25 6.25 -0.07 Salicon [Object] 57.9 76.4 48.0 6.12 0.05 BLIFT[Likes/Views] 58.2 76.2 47.12 6.15 0.38 BLIFT[Titles] 58.4 78.1 48.12 5.09 0.13 BLIFT[Comments] 59.0 79.1 49.58 3.02 0.19 BLIFT 59.2 79.3 50.38 3.05 0.40